Let X1 and X2 be independent n(0,1) random variables. Find the pdf of (X1 - X2)^2/2
Answer #1

TOPIC:Transformation of random variables.

Given that, Xom N (01) X₂ ~ N(OID Yindependently. MER. Now, the pdf of x, is- 67, (d) e So, the mgf of X, is - M4, C4) - E[etSimilarly, X2 ~N(0,1). Sun My, (t) = € (e tk). Define y = x - x - So, the mgf of y is - - My (t) = Elett) Due to independence(5/2/2 ( 4 )/2 e 7. eth/ t/ - =e - which is the mgf of NOOID. By the uniqueness of mas; we get, • Y = (x1-x2) WN(01). 2) DefiSo, the cof of w is - > Fw (W) = p (wsw). =P (y2 300 = P(11 125) = P(VW SY EWW). * P/vw) - (-). So, the pch of w is - Sw (W)

-which is the required pdf.